OpenStudy (anonymous):

I need help in solving this equation: (3x^2+5)/(7x+5)=2

OpenStudy (anonymous):

5,-1/3

OpenStudy (anonymous):

(3x^2 + 5)/(7x + 5) = 2 3x^2 + 5= 2(7x+5) 3x^2 - 14x + 5 - 10 = 0 3x^2 - 14x - 5 = 0 3x^2 -15x + x -5 = 0 3x(x-5) + 1(x-5) = 0 (3x+1)(x-5) = 0 So, (3x+1) = 0 (or) (x-5) = 0 Therefore, x= -1/3 (or) x=5